OpenStudy (llroberts43):

How do I find the equation of a line that passes through (-2,-1) and (-1, -6)? I have to put it in slope intercept form

OpenStudy (calculusxy):

\[\large y = mx+b\] m = slope b = y-intercept

OpenStudy (calculusxy):

First, I suggest you to figure out the slope.

OpenStudy (calculusxy):

Do you know how to?

OpenStudy (llroberts43):

the slope is -5

OpenStudy (calculusxy):

Okay so we will go with that.

OpenStudy (calculusxy):

Now that we have the slope as -5, we can substitute m for that. \[\large y = -5x + b\] Usually, in this case, what I would do is choose a coordinate from the two that are already given. So we will use (-2,-1)... As you can see that with -5 there is a "x". We will substitute that x with -2 and the y before the equal sign will be replaced with the y-coordinate (or -1). \[\large -1 = -5(-2) + b\] Now you need to solve for b.

OpenStudy (llroberts43):

So I'm multiplying -5 and -2 and that gives me 10

OpenStudy (calculusxy):

yes go on...

OpenStudy (llroberts43):

I don't know how to solve for b...I don't know what to do next

OpenStudy (calculusxy):

\[\large -1 = 10 + b\] \[\large -1 - 10 = 10 - 10 + b\] \[\large -11 = b\] Let's check this now: \[\large -1 = -5(-2) - 11 => -1 = 10 - 11 => -1 = -1\]

OpenStudy (calculusxy):

\[\large -6 = -5(-1) -11 => -6 = 5 - 11 => -6 = -6\]

OpenStudy (calculusxy):

So your slope intercept form is \[\large y = -5x - 11\]

OpenStudy (calculusxy):

Do you understand how I got it? @llroberts43

OpenStudy (llroberts43):

sorry I'm at work...trying to get off the phone so I can look

OpenStudy (llroberts43):

Ok, you subtracted 10 from both sides and that gave you the -11

OpenStudy (calculusxy):

yes

OpenStudy (calculusxy):

we needed to isolate b

OpenStudy (llroberts43):

Thank you. I have a final on this stuff today and I'm just stressed out. I don't remember doing this stuff

OpenStudy (calculusxy):

It's okay .. take a deep breath and you will be fine.. just review the work that I did in here.

OpenStudy (llroberts43):

I will. thanks again